Beams Document 1315-v1

DocDB Home ]  [ Search ] [ Last 20 Days ] [ List Authors ] [ List Topics ] [ List Events ]

Performance of New Echotek for the Last Two Shots Before the Shutdown

Document #:
Beams-doc-1315-v1
Document type:
Note
Submitted by:
Robert K Kutschke
Updated by:
Robert K Kutschke
Document Created:
24 Aug 2004, 16:20
Contents Revised:
24 Aug 2004, 16:23
Metadata Revised:
26 Aug 2004, 11:39
Viewable by:
  • Public document
Modifiable by:

Quick Links:
Latest Version

Other Versions:
Abstract:
This note presents data taken using new BPM electronics at HA34 and VA35 for the shots done on August 21, 2004 and August 22, 2004. These are the last two shots before the fall 2004 shutdown. A bug was found in the online cancellation of the proton signal on the anti-proton cables. Otherwise the instrument worked well, as it has throughout the past week.
Files in Document:
Topics:
DocDB Home ]  [ Search ] [ Last 20 Days ] [ List Authors ] [ List Topics ] [ List Events ]

DocDB Version 8.8.9, contact Beams Document Database Administrators